The perimeter of a rectangle is 50ft. The width of a rectangle is 5ft. less than the length. Find the length and the width of the rectangle. P = 2w +2L
Answer by alok(3)

The perimeter of a rectangle= 2w+2l where w is the width & l is the length.
Here, width is length less 5 ft.
or w=l-5
hence replacing w by l-5 we get P= 2(l-5)+2l=50
or 50= 2l -10+2l
or 50=4l-10
or 60=4l
or l=60/4
or l=15
hence length is 15 ft & width is 10 ft